# Incentivize Contribution and Learn Parameters Too: Federated Learning with Strategic Data Owners

### v2

@@ 3 Learning with arbitrary effective costs
In this section, we consider the effective cost of agent $i$, i.e., $z_{i}(s_{i},s_{-i})$, to be an arbitrary function of $s=(s_{i},s_{-i})$. We know that an alternative interpretation of NE (Definition 1) is a strategy profile $(s_{i}^{*},s_{-i}^{*})$ where every agent’s best response to the strategies of the other players is its own strategy in that profile (see  [22, e.g.]), i.e., $s_{i}^{*}\in\operatorname*{\arg\!\max}_{s_{i}\in S_{i}}\ u_{i}(w,s_{i},s_{-i}^{*}),\ \forall i\in N$. Hence, an algorithm that simultaneously updates all agents’ strategies with the best responses to the current strategies of the other players is called a best response dynamics of a strategic form game (see  [7, e.g.] for a detailed description). In our problem, this approach cannot be directly employed since the center also needs to learn and update the model parameters $w$ as the agents choose their data contributions. We, therefore, propose a mechanism for federated learning that simultaneously updates both the agents’ strategies and the center’s choice of $w$.
In this mechanism, each agent $i$ starts with some initial choices of $s_{i}^{0}$ and shares that with the center. [^1] Note that, only the number of data points, $s_{i}$ is shared with the center and not the data, which is consistent with the principle of federated learning. The center also starts with a $w^{0}$ and broadcasts that and the entire initial data-contribution vector $s^{0}$ with all the agents. In every subsequent iteration $t$, each agent $i$ locally computes two quantities: (i) her updated contribution $s_{i}^{t+1}$ by taking one gradient ascent step w.r.t. her own contribution $s_{i}$, and (ii) the local gradient of agent $i$’s accuracy component of the social welfare w.r.t. $w$. Both are evaluated at the current values of $w^{t}$ and $s^{t}$ and sent back to the center. The center calculates an updated $w^{t+1}$ that averages (in spirit of the FedAvg algorithm [23]) all the local gradients sent by the agents. The center then shares $w^{t+1}$ and $s^{t+1}$ with all the agents. Formally, the updates are given as follows.
@@
where the function $\tilde{g}(w^{t},s^{t})=\frac{1}{n}\sum_{i\in N}\nabla_{w}a_{i}(w^{t},s^{t})$ and $g$ is defined as
@@
[…]
The mechanism is detailed out in Algorithm 1. Our main result of this section is that under certain bounded derivative conditions, Algorithm 1 always converges to a Nash equilibrium. We need a few matrices, defined as follows for every $w$ and $s$. For $i,j\in N$ and $k,\ell\in\{1,\ldots,m\}$,
$\begin{split}G(w,s)_{ij}&=\frac{\partial^{2}}{\partial s_{j}\partial s_{i}}u_{i}(w,s);\\ \tilde{G}(w,s)_{k\ell}&=\frac{1}{n}\sum_{i\in N}\frac{\partial^{2}}{\partial w_{\ell}\partial w_{k}}a_{i}(w,s);\\ H(w,s)_{ik}&=\frac{\partial^{2}}{\partial w_{k}\partial s_{i}}u_{i}(w,s);\\ \tilde{H}(w,s)_{kj}&=\sum_{i\in N}\frac{\partial^{2}}{\partial s_{j}\partial w_{k}}a_{i}(w,s).\end{split}$ (5)
@@
Consider the utility functions given by Equation 1 where functions $a_{i},z_{i},i\in N$ are such that the following properties hold for every $w\in\mathbb{R}^{m}$ and $s\in\prod_{i\in N}S_{i}$ (the matrices below are as defined in Equation 5).
@@
    We assume the following bounds: (i) $\forall i,j\in N,|G(w,s)_{ij}|\leqslant L$, (ii) $\forall k,\ell\in\{1,\ldots,m\},|\tilde{G}(w,s)_{k\ell}|\leqslant\tilde{L}$, (iii) $\|H(w,s)\|_{op}\leqslant P$, where $\|A\|_{op}:=\inf\{c\geqslant 0:\|Av\|\leqslant c\|v\|,\forall v\}$, (iv) $\|\tilde{H}(w,s)\|_{op}\leqslant\tilde{P}$.
Note that the definition of $G$ and $\tilde{G}$ do not imply the strong concavity of $u_{i}$ or $a_{i}$. Moreover, the above assumptions are standard and have appeared in the literature before (see [25]). We define the following expressions for a cleaner presentation.
$\begin{split}W_{1}=\sqrt{1+\gamma^{2}n^{2}L^{2}-2\gamma\lambda}&+\sqrt{\tilde{P}^{2}\gamma^{2}},\\ W_{2}=\sqrt{1+\eta^{2}m^{2}\tilde{L}^{2}-2\eta\tilde{\lambda}}&+\sqrt{P^{2}\eta^{2}},\\ E=\|g(w^{0},s^{0},\mu^{0})\|_{2}+\|\tilde{g}(w^{0},s^{0})\|_{2},&\quad\text{ where }w^{0},s^{0}\text{ are arbitrary, and}\\ T_{0}(w^{0},s^{0})=\nicefrac{{\left(\ln{\frac{\displaystyle E}{\displaystyle\epsilon}}\right)}}{{\left(\ln{\frac{\displaystyle 1}{\displaystyle W}}\right)}},&\quad\text{ where }W=\max\{W_{1},W_{2}\}.\end{split}$ (6)
@@
Under 1, UPBReD (Algorithm 1) converges to a Nash equilibrium. Formally, for a given $\epsilon>0$, for every initial value $(w^{0},s^{0})$ of Algorithm 1, the gradients $\|g(w^{T},s^{T},\mu^{T})\|<\epsilon$ and $\|\tilde{g}(w^{T},s^{T})\|<\epsilon$, for all $T\geqslant T_{0}(w^{0},s^{0})$, when the step sizes are chosen as follows:
@@ 4       Center broadcasts $w^{t},s^{t}$
7             $s^{t+1}=s^{t}+\gamma g(w^{t},s^{t},\mu^{t})$
8             Compute local gradient: $d_{i}^{t+1}=\nabla_{w}\>a_{i}(w^{t},s_{i}^{t},s_{-i}^{t})$
@@ 9             Send $s_{i}^{t+1},d_{i}^{t+1}$ to the center
The theorem provides a guarantee of convergence of Algorithm 1 for any arbitrary initial condition $(w^{0},s^{0})$. It needs to run for a minimum number of iterations $T_{0}(w^{0},s^{0})$, with appropriate parameters of $\gamma$ and $\eta$ that govern the gradient ascent rates of the agents’ and center’s objectives respectively. Note that for a better convergence, i.e., a smaller $\epsilon$, the algorithm needs to run longer. The parameter $W$ determines the contraction factor of the recurrence of $\|g(w^{t},s^{t})\|_{2}+\|\tilde{g}(w^{t},s^{t})\|_{2}$ and is chosen to be smaller than unity by choosing $\gamma$ and $\eta$ appropriately. Section A.2 shows an example of such a feasible region.

### v3

@@ 3 Learning with arbitrary effective costs
In this section, we consider the effective cost of agent $i$, i.e., $z_{i}(s_{i},s_{-i})$, to be an arbitrary function of $s=(s_{i},s_{-i})$. We know that an alternative interpretation of NE (Definition˜1) is a strategy profile $(s_{i}^{*},s_{-i}^{*})$ where every agent’s best response to the strategies of the other players is its own strategy in that profile (see  (Maschler et al., 2020, e.g.)), i.e., $s_{i}^{*}\in\operatorname*{\arg\!\max}_{s_{i}\in S_{i}}\ u_{i}(w,s_{i},s_{-i}^{*}),\ \forall i\in N$. Hence, an algorithm that simultaneously updates all agents’ strategies with the best responses to the current strategies of the other players is called a best response dynamics of a strategic form game (see  (Fudenberg, 1991, e.g.) for a detailed description). In our problem, this approach cannot be directly employed since the center also needs to learn and update the model parameters $w$ as the agents choose their data contributions. We, therefore, propose a mechanism for federated learning that simultaneously updates both the agents’ strategies and the center’s choice of $w$. Federated learning protocols iteratively use local gradients to update a global model, the mechanism we describe below captures strategic behaviour at every iteration by allowing agents to update their strategy (data set sizes used for gradient computation) at every realisation of the global model based on Equation˜1
In this mechanism, each agent $i$ starts with some initial choices of $s_{i}^{0}$ and shares that with the center. [^1] Note that, only the number of data points, $s_{i}$ is shared with the center and not the data, which is consistent with the principle of federated learning. The center also starts with a $w^{0}$ and broadcasts that and the entire initial data-contribution vector $s^{0}$ to all the agents. In every subsequent iteration $t$, each agent $i$ locally computes two quantities: (i) its updated contribution $s_{i}^{t+1}$ by taking one gradient ascent step w.r.t. its own contribution $s_{i}$, and (ii) the local gradient of agent $i$’s valuation component of the social welfare w.r.t. $w$. Both are evaluated at the current values of $w^{t}$ and $s^{t}$ and sent back to the center. The center calculates an updated $w^{t+1}$ that averages (in spirit of the FedAvg algorithm (McMahan et al., 2017)) all the local gradients sent by the agents. The center then shares $w^{t+1}$ and $s^{t+1}$ with all the agents. Formally, the updates are given as follows.
@@
where the function $\tilde{g}(w^{t},s^{t})=\frac{1}{n}\sum_{i\in N}\nabla_{w}v_{i}(w^{t},s^{t})$ and $g$ is defined as
@@
$\text{ where }\mu_{i}^{t}$ $=\begin{cases}-\frac{\partial}{\partial s_{i}}u_{i}(w^{t},s^{t}),&\text{when either }s_{i}^{t}=0,\frac{\partial}{\partial s_{i}}u_{i}<0\\ &\text{or }s_{i}^{t}=s_{i}^{\max},\frac{\partial}{\partial s_{i}}u_{i}>0\\ 0.&\text{otherwise}\end{cases}$
The mechanism is detailed out in Algorithm˜1. Our main result of this section is that under certain bounded derivative conditions, Algorithm˜1 always converges to a Nash equilibrium. We need a few matrices, defined as follows for every $w$ and $s$. For $i,j\in N$ and $k,\ell\in\{1,\ldots,m\}$,
$\begin{split}\textstyle&\textstyle G(w,s)_{ij}=\frac{\partial^{2}}{\partial s_{j}\partial s_{i}}u_{i}(w,s);\>\tilde{G}(w,s)_{k\ell}=\frac{1}{n}\sum_{i\in N}\frac{\partial^{2}}{\partial w_{\ell}\partial w_{k}}v_{i}(w,s);\\ \textstyle&\textstyle H(w,s)_{ik}=\frac{\partial^{2}}{\partial w_{k}\partial s_{i}}u_{i}(w,s);\>\tilde{H}(w,s)_{kj}=\sum_{i\in N}\frac{\partial^{2}}{\partial s_{j}\partial w_{k}}v_{i}(w,s).\end{split}$ (5)
@@
Consider the utility functions given by Equation˜1 where functions $v_{i},z_{i},i\in N$ are such that the following properties hold for every $w\in\mathbb{R}^{m}$ and $s\in\prod_{i\in N}S_{i}$ (the matrices below are as defined in Equation˜5).
@@
    We assume the following bounds:
    1. (a)
      $\forall i,j\in N,|G(w,s)_{ij}|\leqslant L$,
    2. (b)
      $\forall k,\ell\in\{1,\ldots,m\},|\tilde{G}(w,s)_{k\ell}|\leqslant\tilde{L}$,
    3. (c)
      $\|H(w,s)\|_{op}\leqslant P$, where $\|A\|_{op}:=\inf\{c\geqslant 0:\|Ax\|\leqslant c\|x\|,\forall x\}$,
    4. (d)
      $\|\tilde{H}(w,s)\|_{op}\leqslant\tilde{P}$.
Note that the definition of $G$ and $\tilde{G}$ do not imply the strong concavity of $u_{i}$ or $v_{i}$. Moreover, the above assumptions are standard and have appeared in the literature before (see (Murhekar et al., 2023)). We define the following expressions for a cleaner presentation.
$\begin{split}\textstyle W_{1}=\sqrt{1+\gamma^{2}n^{2}L^{2}-2\gamma\lambda}&\textstyle+\sqrt{\tilde{P}^{2}\gamma^{2}},\\ \textstyle W_{2}=\sqrt{1+\eta^{2}m^{2}\tilde{L}^{2}-2\eta\tilde{\lambda}}&\textstyle+\sqrt{P^{2}\eta^{2}},\\ \textstyle E=\|g(w^{0},s^{0},\mu^{0})\|_{2}+\|\tilde{g}(w^{0},s^{0})\|_{2},&\textstyle\>\text{ where }w^{0}\in\mathbb{R}^{m},s^{0}\in\prod_{i\in N}S_{i}\\ \textstyle T_{0}(w^{0},s^{0})=\nicefrac{{\left(\ln{\frac{\textstyle E}{\textstyle\epsilon}}\right)}}{{\left(\ln{\frac{\textstyle 1}{\textstyle W}}\right)}},&\textstyle\>\text{ where }W=\max\{W_{1},W_{2}\}.\end{split}$ (6)
@@
Under ˜1, UPBReD (Algorithm˜1) converges to a Nash equilibrium. Formally, for a given $\epsilon>0$, for every initial value $(w^{0},s^{0})$ of Algorithm˜1, the gradients $\|g(w^{T},s^{T},\mu^{T})\|<\epsilon$ and $\|\tilde{g}(w^{T},s^{T})\|<\epsilon$, for all $T\geqslant T_{0}(w^{0},s^{0})$, when the step sizes are chosen as follows:
@@
Proof.
$\quad+{H(w^{\prime},s^{t},\mu^{t})(w^{t+1}-w^{t})},$
$\quad+\tilde{H}(w^{t},s^{\prime})(s^{t+1}-s^{t}),$
Using the update rules of Algorithm˜1 given by Equation˜3 and Equation˜4, we get
$\quad+{H(w^{\prime},s^{t},\mu^{t})\eta\tilde{g}(w^{t},s^{t})}$
$\quad+\tilde{H}(w^{t},s^{\prime})\gamma g(w^{t},s^{t},\mu^{t})$
Using triangle inequality on each of these identities, we get
$\begin{aligned} \textstyle\|g(w^{t+1},s^{t+1},\mu^{t+1})\|_{2}&\textstyle\leqslant\|(I_{n\times n}+\gamma G(w^{t},s^{\prime},\mu^{t}))g(w^{t},s^{t},\mu^{t})\|_{2}\\ \textstyle&\textstyle\quad+\eta\|{H(w^{\prime},s^{t},\mu^{t})\tilde{g}(w^{t},s^{t})}\|_{2}\\ \textstyle\|\tilde{g}(w^{t+1},s^{t+1})\|_{2}&\textstyle\leqslant\|(I_{m\times m}+\eta\tilde{G}(w^{\prime},s^{t}))\tilde{g}(w^{t},s^{t})\|_{2}\\ \textstyle&\textstyle\quad+\gamma\|\tilde{H}(w^{t},s^{\prime})g(w^{t},s^{t},\mu^{t})\|_{2}\end{aligned}$ (7)
From condition 1 of ˜1, we get that $v^{\top}(G+\lambda I_{n\times n})v\leqslant 0,\forall v\in\mathbb{R}^{n}$ and $v^{\prime\top}(\tilde{G}+\lambda I_{m\times m})v^{\prime}\leqslant 0,\forall v^{\prime}\in\mathbb{R}^{m}$. In particular, for $v=g(w^{t},s^{t},\mu^{t})$ and $v^{\prime}=\tilde{g}(w^{t},s^{t})$, we get
$\begin{split}\textstyle g(w^{t},s^{t},\mu^{t})^{\top}G(w^{t},s^{\prime},\mu^{t})g(w^{t},s^{t},\mu^{t})&\textstyle\leqslant-\lambda\|g(w^{t},s^{t},\mu^{t})\|_{2}^{2}\\ \textstyle\tilde{g}(w^{t},s^{t})^{\top}\tilde{G}(w^{\prime},s^{t})\tilde{g}(w^{t},s^{t})&\textstyle\leqslant-\tilde{\lambda}\|\tilde{g}(w^{t},s^{t})\|_{2}^{2}\end{split}$ (8)
Consider the square of the first term of the RHS of the inequality of $g$ in Equation˜7
[…]
where the equality comes by expanding the squared norm and the inequality comes from the facts that (i) $\|G(w^{t},s^{t})g(w^{t},s^{t},\mu^{t})\|_{2}^{2}\leqslant\|G(w^{t},s^{t})\|_{F}^{2}\|g(w^{t},s^{t},\mu^{t})\|_{2}^{2}$, where $\|A\|_{F}:=\sqrt{\sum_{i}\sum_{j}|A_{ij}|^{2}}$ is the Frobenius norm of a matrix $A$ and (ii) using Equation˜8. By condition 2 of ˜1, $\|G(w^{t},s^{t})\|_{F}^{2}\leqslant n^{2}L^{2}$. Hence, we get
$\textstyle\|(I_{n\times n}+\gamma.G(w^{t},s^{t}))g(w^{t},s^{t},\mu^{t})\|$
$\textstyle\leqslant\sqrt{1+\gamma^{2}.n^{2}L^{2}-2\gamma\lambda}\cdot\|g(w^{t},s^{t},\mu^{t})\|_{2}$
given the term inside the square root is positive.
Consider the second term of the RHS of the inequality of $g$ in Equation˜7, where
$\textstyle\|H(w^{\prime},s^{t},\mu^{t})\tilde{g}(w^{t},s^{t})\|_{2}^{2}$ $\textstyle\leqslant\|H(w^{\prime},s^{t},\mu^{t})\|_{op}^{2}\|\tilde{g}(w^{t},s^{t})\|_{2}^{2}$
$\textstyle\leqslant P^{2}\|\tilde{g}(w^{t},s^{t})\|_{2}^{2}$
by condition 2 of ˜1.
Defining $\alpha:=1+\gamma^{2}.n^{2}L^{2}-2\gamma\lambda,\beta:=P^{2}\eta^{2},\tilde{\alpha}:=1+\eta^{2}.m^{2}\tilde{L}^{2}-2\eta\tilde{\lambda},\tilde{\beta}:=\tilde{P}^{2}\gamma^{2}$, and carrying out a similar analysis for $\tilde{g}$ in Equation˜7, we get
$\begin{split}\textstyle\|g(w^{t+1},s^{t+1},\mu^{t+1})\|_{2}&\textstyle\leqslant\sqrt{\alpha}\|g(w^{t},s^{t},\mu^{t})\|_{2}+\sqrt{\beta}\|\tilde{g}(w^{t},s^{t})\|_{2}\\ \textstyle\|\tilde{g}(w^{t+1},s^{t+1})\|_{2}&\textstyle\leqslant\sqrt{\tilde{\alpha}}\|\tilde{g}(w^{t},s^{t})\|_{2}+\sqrt{\tilde{\beta}}\|g(w^{t},s^{t},\mu^{t})\|_{2}\end{split}$ (10)
Adding the inequalities of Equation˜10
[…]
To ensure that the above inequality is a contraction, we need to ensure $\sqrt{\alpha}+\sqrt{\tilde{\beta}},\sqrt{\tilde{\alpha}}+\sqrt{\beta}\in(0,1)$. These imply (i) $0<\alpha<1,0<\tilde{\alpha}<1$, (ii) $\sqrt{\tilde{\beta}}<1,\sqrt{\beta}<1$, and (iii) $\sqrt{\alpha}+\sqrt{\tilde{\beta}}<1,\sqrt{\tilde{\alpha}}+\sqrt{\beta}<1$. We can solve for $\gamma$ and $\eta$ from these inequalities and obtain the sufficient conditions when $\lambda>\tilde{P}$ and $\tilde{\lambda}>P$:
Notice that from conditions 1 and 2 of ˜1, we have that $\lambda<L$ and $\tilde{\lambda}<\tilde{L}$. Thus $n^{2}L^{2}-\tilde{P}^{2}>0$ and $m^{2}\tilde{L}^{2}-P^{2}>0$. For the $\gamma,\eta$ chosen as above, we find $W_{1}=\sqrt{\alpha}+\sqrt{\tilde{\beta}}<1$ and $W_{2}=\sqrt{\tilde{\alpha}}+\sqrt{\beta}<1$, and hence $W=\max\{W_{1},W_{2}\}<1$. Therefore,
$\|g(w^{t+1},s^{t+1},\mu^{t+1})\|_{2}+\|\tilde{g}(w^{t+1},s^{t+1})\|_{2}$
Recursively iterating over this inequality, we get
$\textstyle\|g(w^{T},s^{T},\mu^{T})\|_{2}+\|\tilde{g}(w^{T},s^{T})\|_{2}$
$\textstyle\leqslant W^{T}\left(\|{g}(w^{0},s^{0},\mu^{0})\|_{2}+\|\tilde{g}(w^{0},s^{0})\|_{2}\right).$
Defining $E=\|g(w^{0},s^{0},\mu^{0})\|_{2}+\|\tilde{g}(w^{0},s^{0})\|_{2}$ and
$T_{0}(w^{0},s^{0})=\nicefrac{{\left(\ln{\frac{\textstyle E}{\textstyle\epsilon}}\right)}}{{\left(\ln{\frac{\textstyle 1}{\textstyle W}}\right)}},$
we get that for all $T\geqslant T_{0}(w^{0},s^{0})$,
$\|g(w^{T},s^{T},\mu^{T})\|_{2}+\|\tilde{g}(w^{T},s^{T})\|_{2}<\epsilon$
This completes the proof.
@@ 4 Center broadcasts $w^{t},s^{t}$
7    $s^{t+1}_{i}=s^{t}_{i}+\gamma[g(w^{t},s^{t},\mu^{t})]_{i}$
8      Compute local gradient: $d_{i}^{t+1}=\nabla_{w}\>v_{i}(w^{t},s_{i}^{t},s_{-i}^{t})$
@@ 9      Send $s_{i}^{t+1},d_{i}^{t+1}$ to the center
The theorem provides a guarantee of convergence of Algorithm˜1 for any arbitrary initial condition $(w^{0},s^{0})$. It needs to run for a minimum number of iterations $T_{0}(w^{0},s^{0})$, with appropriate parameters of $\gamma$ and $\eta$ that govern the gradient ascent rates of the agents’ and center’s objectives respectively. Note that for a better convergence, i.e., a smaller $\epsilon$, the algorithm needs to run longer. The parameter $W$ determines the contraction factor of the recurrence of $\|g(w^{t},s^{t})\|_{2}+\|\tilde{g}(w^{t},s^{t})\|_{2}$ and is chosen to be smaller than unity by choosing $\gamma$ and $\eta$ appropriately. The following example shows an example of such a feasible region.
Example 1 (Feasible region of $\gamma$ for Theorem˜1 to hold).
Figure˜1 shows feasible regions of $\gamma$ for certain $n,L,\lambda,\tilde{P}$.
(a) $\tilde{P}=1$
(b) $\lambda=1.1$